OpenStudy (anonymous):

FIND dy/dx by implicit differentiation: x^3+y^3=8

OpenStudy (anonymous):

3x² +3y²y' = 0 y' = -x²/y²

OpenStudy (anonymous):

but y = (8 - x³)^(1/3) so y' = -x²/(8-x²)^(2/3)

OpenStudy (anonymous):

@JoãoVitorMC thank you (:

OpenStudy (anonymous):

welcome xD